OpenStudy (anonymous):

Find tan θ if cot θ = 2 and 0° < θ < 90°.

OpenStudy (anonymous):

\[\bf \cot(\theta)=\frac{ 1 }{ \tan(\theta) }\]\[\bf \frac{ 1 }{ \tan(\theta) }=\cot(\theta)=2 \implies \tan(\theta) =\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Thank you :)